# The Andoni-Naor-Neiman inequalities and isometric embeddability into a $\mathrm{CAT}(0)$ space

### v1

@@ 3. Lebedeva’s 6-point metric space
For any two points $a$ and $b$ in the Euclidean space $\mathbb{R}^{3}$, we denote by $[a,b]$ the line segment joining $a$ and $b$, and by $(a,b)$ the set $[a,b]\setminus\{a,b\}$. We also denote by $\mathrm{conv}(S)$ the convexhull of a subset $S$ of $\mathbb{R}^{3}$. Suppose $x_{1},\ldots,x_{6}$ are distinct six points in $\mathbb{R}^{3}$ that satisfy the following conditions:
@@
$H$ $=\max_{y\in\mathrm{conv}\left(L\setminus\left\{x_{6}\right\}\right)}\|x_{6}-y\|,\quad\theta=\min_{k\in[4]}\measuredangle x_{k}x_{6}x_{5},$
where we denote by $\measuredangle x_{k}x_{6}x_{5}\in[0,\pi]$ the interior angle measure at $x_{6}$ of the triangle with vertices $x_{k}$, $x_{6}$ and $x_{5}$. It follows from (7) and (8) that $h>0$, $H>0$ and $\theta>0$. It also follows from (7) and (8) that we can choose $\delta\in(0,\infty)$ such that
@@
holds true for every $\varepsilon>0$ by definition of $d_{\varepsilon}$. So henceforth we assume that $\pi_{56}>0$ and $\pi_{65}>0$. We set
$t_{i}=\frac{\pi_{i6}}{\pi_{i6}+\pi_{6i}}.$
for each $i\in[5]$ with $\pi_{i6}+\pi_{6i}>0$.
@@
$\{z_{p},z_{q}\}\subseteq B\left(x_{5},\frac{h}{2}\right)\cup B\left(x_{6},\frac{h}{2}\right),$
which implies that
$\|z_{p}-z_{ij}\|>\frac{h}{2},\quad\|z_{q}-z_{ij}\|>\frac{h}{2}$ (14)
for any $i,j\in[4]$ with $\pi_{ij}+\pi_{ji}>0$.
We define $K_{0}\in[0,\infty)$ to be the nonnegative real number that satisfies
@@
$\|z-z_{ij}\|^{2}>\frac{h^{2}}{4}\geq\frac{h^{2}}{4}t_{5}$ (19)
for any $i,j\in[4]$. Set

### v2

@@ 3. Lebedeva’s 6-point metric space
For any two points $a$ and $b$ in the Euclidean space $\mathbb{R}^{3}$, we denote by $[a,b]$ the line segment joining $a$ and $b$, and by $(a,b)$ the set $[a,b]\setminus\{a,b\}$. We denote by $\mathrm{conv}(S)$ the convexhull of a subset $S$ of $\mathbb{R}^{3}$. Suppose $x_{1},\ldots,x_{6}$ are distinct six points in $\mathbb{R}^{3}$ that satisfy the following conditions:
@@
$H$ $=\max\left\{\max_{y\in\mathrm{conv}\left(L\setminus\left\{x_{5}\right\}\right)}\|x_{5}-y\|,\quad\max_{y\in\mathrm{conv}\left(L\setminus\left\{x_{6}\right\}\right)}\|x_{6}-y\|\right\},$
$\theta$ $=\min\left\{\min_{k\in[4]}\measuredangle x_{k}x_{5}x_{6},\quad\min_{k\in[4]}\measuredangle x_{k}x_{6}x_{5}\right\}$
where we denote by $\measuredangle x_{k}x_{i}x_{j}\in[0,\pi]$ the interior angle measure at $x_{i}$ of the triangle with vertices $x_{k}$, $x_{i}$ and $x_{j}$. It follows from (7) and (8) that $h>0$, $H>0$ and $\theta>0$. It also follows from (7) and (8) that we can choose $\delta\in(0,h/2)$ such that we have
@@
holds true for every $\varepsilon>0$ by definition of $d_{\varepsilon}$. So henceforth we assume that $\pi_{56}>0$ and $\pi_{65}>0$.
@@
$\{z_{p},z_{q}\}\subseteq B\left(x_{5},\frac{h}{2}\right)$ (14)
or
$\{z_{p},z_{q}\}\subseteq B\left(x_{6},\frac{h}{2}\right).$ (15)
Because it is easily seen that we can obtain the same estimate of $F(\varepsilon)$ in the same way as we describe below whether (14) holds or (15) holds, we assume that (15) holds.
We set
$t_{i}=\frac{\pi_{i6}}{\pi_{i6}+\pi_{6i}}.$
for each $i\in[5]$ with $\pi_{i6}+\pi_{6i}>0$, and define $K_{0}\in[0,\infty)$ to be the nonnegative real number that satisfies
@@
$\|z_{p}-z_{ij}\|>\frac{h}{2},\quad\|z_{q}-z_{ij}\|>\frac{h}{2}$
for any $i\in[4]$ and any $j\in[5]$. Therefore, we have
$\pi_{ij}\|z-z_{ij}\|^{2}>\frac{h^{2}}{4}\pi_{ij}\geq\frac{h^{2}}{4}\pi_{ij}t_{5}$ (20)
for any $i\in[4]$ and any $j\in[5]$. Set